# Spot the Fake: Large Multimodal Model-Based Synthetic Image Detection with Artifact Explanation

### v1

@@ 4.1 Re-thinking LMMs’ Challenges in Synthetic Image Detection
Recent studies have shown that directly using LMMs for synthetic image detection still faces significant challenges [58, 28, 18]. Although large models possess strong text explanation capabilities, when tasked with determining whether an image was AI-generated or identifying forged images from a set, pretrained LMMs often fail to achieve satisfactory performance. This phenomenon highlights the difficulty of relying on LMMs for authenticity judgment, which is closely related to the fact that these models are not inherently designed for synthetic data detection tasks. Nevertheless, through extensive pretraining tasks, multimodal large models have developed strong visual feature extraction abilities and alignment with text. This raises the question: do the internal representations of these large models potentially encode information that can distinguish real images from synthetic ones?
Inspired by Zhang et al. [62], we explored a simple yet effective method on FakeClue: extracting visual features from the last layer of a pre-trained LMM and training a lightweight linear classifier to determine image authenticity. If the representations learned by the model indeed contain discriminative information related to authenticity, even a simple classifier can perform initial detection using these features. As shown in Figure 3, the results confirm that the LMM has potential in distinguishing authenticity. However, in contrast to the conclusions of Zhang et al. [62], framing the task as ”Does the image look real/fake?” by using fixed answers like ”Real” or ”Fake” not only limits the model’s ability to provide textual explanations but also results in suboptimal performance. This is likely due to large models’ challenges in aligning complex visual content with such binary answers. Building on the dataset we constructed, we found that framing the task as visual question answering, where the model is required to provide not just “Real”/“Fake” answers but also explain the image “artifacts”, leads to better alignment between the response text and the image content in Figure 3 (For detailed results, please refer to Appendix). This approach not only improves the performance of artifact explanation but also significantly enhances the overall performance of synthetic image detection.
@@ 4.2 Model Architecture
Global Image Encoder: We employ the pretrained vision backbone of CLIP-ViT(L-14) [35] as our global image encoder. The encoder processes input images with a resolution of 336×336 to preserve synthetic artifact details, resulting in 576 patches per image.
$V=\text{CLIP-ViT}(I)\in\mathbb{R}^{N\times d_{v}}$ (2)
@@
where $W_{1}\in\mathbb{R}^{1024\times 4096}$, $W_{2}\in\mathbb{R}^{4096\times 4096}$ are learnable parameters. The projected features $Z\in\mathbb{R}^{N\times 4096}$ combine with text embeddings of the task prompt $P$ through concatenation.
@@
Method FakeClue (Ours) LOKI Evaluations (ICLR 2025) Acc $\uparrow$ F1 $\uparrow$ ROUGE_L $\uparrow$ CSS $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ ROUGE_L $\uparrow$ CSS $\uparrow$ Deepseek-VL2-small [50] $0.404$ $0.542$ $0.171$ $0.504$ $0.253$ $0.387$ $0.164$ $0.391$ Deepseek-VL2 [50] $0.475$ $0.541$ $0.172$ $0.505$ $0.431$ $0.392$ $0.169$ $0.388$ InternVL2-8B [7] $0.506$ $0.490$ $0.180$ $0.581$ $0.526$ $0.340$ $0.179$ $0.472$ InternVL2-40B [7] $0.507$ $0.463$ $0.176$ $0.552$ $0.507$ $0.376$ $0.184$ $0.473$ Qwen2-VL-7B [46] $0.457$ $0.592$ $0.266$ $0.565$ $0.571$ $0.350$ $0.182$ $0.384$ Qwen2-VL-72B [46] $0.578$ $0.565$ $0.175$ $0.544$ $0.554$ $0.409$ $0.173$ $0.432$ GPT-4o [33] $0.474$ $0.420$ $0.134$ $0.407$ 0.634 0.572 0.147 $0.354$ FakeVLM $\mathbf{0.986}$ $\mathbf{0.981}$ $\mathbf{0.580}$ $\mathbf{0.877}$ $\mathbf{0.843}$ $\mathbf{0.837}$ $\mathbf{0.201}$ $\mathbf{0.582}$
Table 2: The experimental results on the FakeClue and LOKI datasets include both Detection and Artifact Explanation performance.
$\mathcal{L}(\theta)=-\sum_{t=1}^{T_{i}}\log p_{\theta}\left(a_{i,t}\mid a_{i,<t},[Z;E(P)]\right)$ (4)
where $E(\cdot)$ denotes text embeddings. we update all parameters $\theta$ of the LLM during training, enabling comprehensive adaptation to synthetic data reasoning while maintaining the model’s original instruction-following capabilities through full-parameter optimization.
@@ 4.3 Training strategy
As described in the construction process of FakeClue, we leverage category knowledge and utilize multiple LMMs to annotate image artifacts in natural language. As shown on the left side of Fig. 4, we obtain high-quality artifact descriptions as target outputs for the model. Then, we use the QA pairs obtained after the multi-model annotation and summarization steps as our training data. Each data sample consists of: (1) an image $I$; (2) a standardized prompt $P$: ”Does the image look real/fake?”; (3) the aggregated answer $A$ from the multi-model annotations.
Our model was initialized from the original weights of the LLaVA-1.5 7B model [29], and we performed full-parameter fine-tuning using the QA pairs from our constructed dataset. The training is conducted for two epochs on eight NVIDIA A100 GPUs with a batch size of 32 per GPU. We employ a learning rate of 2e-4, with a linear learning rate warmup over the first 3% of the training steps, followed by cosine decay. By fine-tuning all parameters of the model, we enable comprehensive adaptation to the nuances of synthetic data detection and explanation while preserving its general instruction-following capabilities.
@@ 5 Experiment
Figure 5: Comparative case analysis of synthetic image detection, covering animals, people, objects, documents, and remote sensing. FakeVLM outperforms GPT in precision, comprehensiveness, and relevance, indicating its superior detection and interpretation capabilities.
@@ 5.1 Other Benchmarks
LOKI [58] is a recently proposed benchmark for evaluating multimodal large models in general synthetic detection tasks. Beyond just distinguishing real from fake, LOKI also includes human manually annotated fine-grained image artifacts, enabling a thorough exploration of the model’s ability to explain image artifacts.
FF++ [37] is a widely used benchmark dataset for facial forgery detection, containing face images and videos generated by different types of forgery techniques. The dataset includes forged data created using four common forgery methods: DeepFakes, Face2Face, FaceSwap, and NeuralTextures. We used the commonly employed C23 versions.
DD-VQA [61] is a new dataset for artifact explanation in the face domain, designed to leverage human common-sense perception to assess image authenticity. It includes artifacts like blurred hairlines, mismatched eyebrows, rigid pupils, and unnatural skin shadows. Built on FF++ image data, DD-VQA uses manual artifact annotations in a Visual Question Answering (VQA) format, where the model answers questions about artifacts using common-sense reasoning.
@@
CNNSpot [48] 87.8 88.4 28.4 44.2 40.6 43.3
Gram-Net [30] 22.8 34.1 78.8 88.1 67.4 79.4
Fusing [20] 87.7 86.1 15.5 27.2 40.4 36.5
LNP [3] 63.1 67.4 56.9 72.5 58.2 68.3
UnivFD [32] 89.4 88.3 44.9 61.2 53.9 60.7
AntifakePrompt [4] 91.3 92.5 89.3 91.2 90.6 91.2
SIDA [16] 92.9 93.1 90.7 91.0 91.8 92.4
@@
Table 3: Comparison with other detection methods on the DMimage [10] dataset, using the original weights for each method.
5.2 Experimental setup
Task Settings. The LOKI dataset is used solely as the evaluation set, while training is conducted on the FakeClue dataset, with testing performed on LOKI. For the FF++ and DD-VQA datasets, we use their default training-test splits for evaluation. The evaluation metrics are divided into two tasks: detection and artifact explanation. Classification accuracy is represented by accuracy (Acc) and F1 scores, while artifact explanation accuracy is measured using CSS and ROUGE_L. When evaluating large models, since the results are based on generated text rather than binary classification scores, we follow the DD-VQA [61] task and report ACC instead of AUC. However, for evaluating expert models, we still rely on the AUC metric.
Compared Baselines. For tasks that require both synthetic detection and artifact explanation, such as FakeClue, DD-VQA, and LOKI, we compared various general-purpose LMMs, including advanced closed-source models like GPT-4 and open-source models such as Qwen2-VL, LLaVA, InternVL2, and Deepseek-VL2. We also included the Common-DF method proposed in the DD-VQA dataset paper. For pure synthetic data detection tasks, we further compared with recent state-of-the-art expert detection methods.
5.3 Universal synthetic detection
Table 2 presents a comparison of our method (FakeVLM) with other leading general-purpose large models. It is clear that FakeVLM outperforms on multiple metrics, both in terms of synthetic detection performance and artifact explanation capabilities. Specifically, compared to the current powerful open-source model Qwen2-VL-72B, FakeVLM achieves an average improvement of 36.1% in Acc and 41.3% in F1 on both FakeClue and LOKI. Additionally, LOKI includes an extra evaluation metric for human performance, with an Acc of 80.1, whereas FakeVLM achieves an Acc of 84.30, surpassing human performance. This is likely attributed to FakeVLM’s ability to capture deep, image-level features that are imperceptible to the human eye for accurate authenticity judgment.
We present the qualitative evaluation results of FakeVLM in Figure 5. It is evident that FakeVLM not only accurately identifies potential issues arising from the synthesis process, such as visual artifacts, texture distortions, and structural anomalies, but also provides detailed explanations in natural language. This sets FakeVLM apart from traditional synthetic detection methods, which typically rely on probability thresholds to determine if an image is synthetic. FakeVLM offers intuitive and easily understandable descriptions for the detection results, thereby significantly enhancing the interpretability of the synthetic detection process. Moreover, it helps users make confident decisions when assessing synthetic content and provides reliable grounds.
Method DD-VQA (ECCV 2024) Acc $\uparrow$ F1 $\uparrow$ ROUGE_L $\uparrow$ CSS $\uparrow$ InternVL2-8B [7] 0.569 0.531 0.143 0.511 InternVL2-40B [7] 0.525 0.577 0.222 0.545 Qwen2-VL-7B [46] 0.457 0.589 0.266 0.565 Qwen2-VL-72B [46] 0.595 0.579 0.205 0.566 GPT-4o [33] 0.532 0.317 0.130 0.427 Common-DF-T [13] 0.837 0.876 0.577 - Common-DF-I [13] 0.849 0.884 0.588 - Common-DF-TI [13] 0.875 0.901 0.609 - FakeVLM 0.932 0.931 0.704 0.866
@@ 5.3 Universal synthetic detection
Table 4: The experimental results evaluated on the DD-VQA datasets. Common-DF-T and Common-DF-I represent text or image contrastive losses, respectively, while Common-DF-TI denotes both text and image contrastive losses.
We also present the generalization experiment results of FakeVLM on DMimgae [10] in Table 3, following Huang et al. [16]. The experimental results show that the performance gap between FakeVLM and other expert models is not significant, with FakeVLM even outperforming some of them. FakeVLM does not rely on additional classifiers or expert models, yet it achieves performance comparable to or even exceeding that of expert classifiers while retaining its language capability for artifact explanation. This demonstrates the potential of large models in synthetic detection.
Figure 6: Typical cases on DD-VQA dataset. Our model accurately identifies and explains the synthetic artifacts, demonstrating its effectiveness in fine-grained DeepFake detection and interpretation.
@@
FF-DF FF-F2F FF-FS FF-NT Average
FWA [47] 92.1 90.0 88.4 81.2 87.7
Face X-ray [25] 97.9 98.7 98.7 92.9 95.9
SRM [23] 97.3 97.0 97.4 93.0 95.8
CDFA [53] 99.9 86.9 93.3 80.7 90.2
@@
Table 5: Experimental results of fine-tuning BLIP on DD-VQA dataset, which is built on FF++, including both deepfake detection and answer generation performance.
5.4 DeepFake detection
We next evaluate the performance of FakeVLM on DeepFake detection, and Table 4 shows its results on DD-VQA. FakeVLM not only outperforms general-purpose multimodal large models but also significantly exceeds the performance of the specialized vision-language model, Common-DF. Compared to Common-DF, FakeVLM improves Acc by 5.7%, F1 by 3%, and ROUGE_L, which represents artifact explanation performance, by 9.5%. Figure 5 presents the qualitative results on DD-VQA. FakeVLM is capable of not only describing the artifacts across the entire image but also answering questions based on specific detail regions.
Table 5 presents the evaluation results of FakeVLM on the DeepFake synthesis detection dataset FF++, including its performance across multiple sub-categories such as DeepFakes, Face2Face, FaceSwap, and NeuralTextures. The experimental results demonstrate that FakeVLM continues to exhibit strong performance in these tasks. Not only does it maintain a leading position in overall detection, but its performance across each category is also relatively balanced, avoiding the issue of overfitting.

### v2

@@ 4.1 Re-thinking LMMs’ Challenges in Synthetic Image Detection
Recent studies have shown that directly using LMMs for synthetic image detection still face significant challenges ye2024loki ; li2024fakebench ; jia2024can . Although large models possess strong text explanation capabilities, when tasked with determining whether an image was AI-generated or identifying forged images from a set, pretrained LMMs often fail to achieve satisfactory performance. This phenomenon highlights the difficulty of relying on LMMs for authenticity judgment, which is closely related to the fact that these models are not inherently designed for synthetic data detection tasks. Nevertheless, through extensive pretraining tasks, multimodal large models have developed strong visual feature extraction abilities and alignment with text. This raises the question: do the internal representations of these large models potentially encode information that can distinguish real images from synthetic ones?
Inspired by Zhang et al. zhang2024visually , we explored a simple yet effective method on FakeClue: extracting visual features from the last layer of a pre-trained LMM and training a lightweight linear classifier to determine image authenticity. If the representations learned by the model indeed contain discriminative information related to authenticity, even a simple classifier can perform initial detection using these features. As shown in Figure 3, the results confirm that the LMM has potential in distinguishing authenticity. However, in contrast to the conclusions of Zhang et al. zhang2024visually , framing the task as "Does the image look real/fake?" by using fixed answers like "Real" or "Fake" not only limits the model’s ability to provide textual explanations but also results in suboptimal performance. This is likely due to large models’ challenges in aligning complex visual content with such binary answers. Building on the dataset we constructed, we found that framing the task as visual question answering, where the model is required to provide not just “Real”/“Fake” answers but also explain the image “artifacts”, leads to better alignment between the response text and the image content in Figure 3 (For detailed results, please refer to Appendix B). This approach not only improves the performance of artifact explanation but also significantly enhances the overall performance of synthetic image detection.
@@ 4.2 Model Architecture
Global Image Encoder: We employ the pretrained vision backbone of CLIP-ViT(L-14) clip as our global image encoder. The encoder processes input images with a resolution of 336×336 to preserve synthetic artifact details, resulting in 576 patches per image.
[…]
@@
[…]
@@
[…]
where $E(\cdot)$ denotes text embeddings. We update all parameters of the LLM during training, enabling full adaptation to synthetic data reasoning while preserving original instruction-following capabilities via full-parameter optimization
@@ 4.3 Training strategy
As described in the construction process of FakeClue, we leverage category knowledge and utilize multiple LMMs to annotate image artifacts in natural language. As shown on the left side of Fig. 2, we obtain high-quality artifact descriptions as target outputs for the model. Then, we use the QA pairs obtained after the multi-model annotation and summarization steps as our training data. Each data sample consists of: (1) an image $I$; (2) a standardized prompt $P$: "Does the image look real/fake?"; (3) the aggregated answer $A$ from the multi-model annotations.
Our model, initialized from LLaVA-1.5 7B weights liu2023llava , underwent full-parameter fine-tuning on our constructed QA dataset. The training is conducted for two epochs on eight NVIDIA A100 GPUs with a batch size of 32 per GPU using a 2e-5 learning rate with 3% linear warmup and cosine decay. This full fine-tuning adapted the model to synthetic data detection/explanation nuances while preserving its general instruction-following capabilities.
Table 2: The experimental results on the FakeClue and LOKI datasets include both Detection and Artifact Explanation performance. * denotes methods trained on FakeClue.
Method FakeClue (Ours) LOKI Evaluations (ICLR 2025) Acc $\uparrow\uparrow$ F1 $\uparrow\uparrow$ ROUGE_L $\uparrow\uparrow$ CSS $\uparrow\uparrow$ Acc $\uparrow\uparrow$ F1 $\uparrow\uparrow$ ROUGE_L $\uparrow\uparrow$ CSS $\uparrow\uparrow$ Deepseek-VL2-small wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels $40.4$ $54.2$ $17.1$ $50.4$ $25.3$ $38.7$ $16.4$ $39.1$ Deepseek-VL2 wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels $47.5$ $54.1$ $17.2$ $50.5$ $43.1$ $39.2$ $16.9$ $38.8$ InternVL2-8B chen2024internvl $50.6$ $49.0$ $18.0$ $58.1$ $52.6$ $34.0$ $17.9$ $47.2$ InternVL2-40B chen2024internvl $50.7$ $46.3$ $17.6$ $55.2$ $50.7$ $37.6$ $18.4$ $47.3$ Qwen2-VL-7B wang2024qwen2 $45.7$ $59.2$ $26.6$ $56.5$ $57.1$ $35.0$ $18.2$ $38.4$ Qwen2-VL-72B wang2024qwen2 $57.8$ $56.5$ $17.5$ $54.4$ $55.4$ $40.9$ $17.3$ $43.2$ GPT-4o (2024-08-06) gpt4o $47.4$ $42.0$ $13.4$ $40.7$ $63.4$ $57.2$ $14.7$ $35.4$ CNNSpot wang2020cnn $43.1$ $9.8$ - - $43.1$ $11.4$ - - FreqNet tan2024frequency $48.7$ $39.3$ - - $58.9$ $50.6$ - - Fatformer liu2024forgery $54.5$ $45.1$ - - $58.8$ $48.4$ - - UnivFD ojha2023towards $63.1$ $46.8$ - - $49.0$ $35.8$ - - AIDE yan2024sanity $85.9$ $94.5$ - - $65.6$ $80.2$ - - NPR tan2024rethinking $90.2$ $91.6$ - - $77.4$ $80.0$ - - FakeVLM $\mathbf{98.6}$ $\mathbf{98.1}$ $\mathbf{58.0}$ $\mathbf{87.7}$ $\mathbf{84.3}$ $\mathbf{83.7}$ $\mathbf{20.1}$ $\mathbf{58.2}$
@@ 5.1 Other Benchmarks
LOKI ye2024loki is a recently proposed benchmark for evaluating multimodal large models in general synthetic detection tasks. Beyond just distinguishing real from fake, LOKI also includes human manually annotated fine-grained image artifacts, enabling a thorough exploration of the model’s ability to explain image artifacts. This inclusion allows us to verify, to some extent, whether our model’s perception of artifacts aligns with human cognition.
FF++ roessler2019faceforensicspp is a widely used benchmark dataset for facial forgery detection, containing face images and videos generated by different types of forgery techniques. The dataset includes forged data created using four common forgery methods: DeepFakes, Face2Face, FaceSwap, and NeuralTextures. We used the commonly employed C23 versions.
DD-VQA zhang2024common is a new face-domain artifact explanation dataset leveraging human common-sense perception for authenticity assessment. It features artifacts like blurred hairlines and unnatural skin shadows. Built on FF++ data, DD-VQA employs manual artifact annotations in a VQA format, requiring models to answer common-sense questions about artifacts.
5.2 Experimental setup
Task Settings. LOKI serves as the evaluation set, while training is conducted on the FakeClue dataset, with testing performed on LOKI. For FF++ and DD-VQA, we use their default training-test splits for evaluation. Evaluation metrics cover two tasks: detection and artifact explanation. Classification accuracy is represented by Acc, Auc, and F1 scores, while artifact explanation accuracy is measured using CSS and ROUGE_L.
Compared Baselines. For tasks requiring both synthetic detection and artifact explanation (e.g., FakeClue, DD-VQA, LOKI), we compared various general-purpose LMMs, including closed-source models like GPT-4 and open-source ones such as Qwen2-VL, LLaVA, InternVL2, and Deepseek-VL2. We also included the Common-DF method from the DD-VQA dataset. For pure synthetic data detection, we further compared with recent SOTA expert methods.
5.3 Universal synthetic detection
Table 2 compares FakeVLM with leading general-purpose LMMs and expert models, demonstrating its superior performance in both synthetic detection and artifact explanation. Specifically, compared to the current powerful open-source model Qwen2-VL-72B and leading expert model NPR or AIDE, which is also trained on FakeClue, FakeVLM achieves an average improvement of 7.7% in Acc and 3.6% in F1 on both FakeClue and LOKI. Additionally, LOKI includes an extra evaluation metric for human performance, with an Acc of 80.1%, whereas FakeVLM achieves an Acc of 84.3%, surpassing human performance. This is likely attributed to FakeVLM’s ability to capture deep, image-level features that are imperceptible to the human eye for accurate authenticity judgment. Moreover, FakeVLM’s strong ROUGE_L and CSS scores on the human-annotated LOKI dataset highlight alignment with human artifact perception, a noteworthy achievement given its training on entirely LMM-generated data, which is largely thanks to our meticulously designed data annotation pipeline.
Figure 4 presents FakeVLM’s qualitative evaluation. FakeVLM identifies synthesis issues (e.g., visual artifacts, texture distortions, structural anomalies) with detailed natural language explanations. Unlike probability-threshold methods, FakeVLM’s intuitive, interpretable descriptions, enhance detection transparency, enabling confident and reliable synthetic content assessment.
Figure 4: Synthetic image detection cases, covering animals, people, objects, documents, and remote sensing (red denotes incorrect, green denotes correct detection). FakeVLM outperforms GPT in precision, comprehensiveness, and relevance, demonstrating superior detection and interpretation.
We also present the generalization experiment results of FakeVLM on DMimage corvi2023detection in Table 3, following Huang et al. huang2024sida . The experimental results show that the performance gap between FakeVLM and other expert models is not significant, with FakeVLM even outperforming some of them. FakeVLM does not rely on additional classifiers or expert models, yet it achieves performance comparable to or even exceeding that of expert classifiers while retaining its language capability for artifact explanation. This demonstrates the potential of large models in synthetic detection.
Table 3: Comparison with other detection methods on the DMimage corvi2023detection dataset, using the original weights for each method.
@@ 5.3 Universal synthetic detection
CNNSpot wang2020cnn 87.8 88.4 28.4 44.2 40.6 43.3
Gram-Net liu2020global 62.8 54.1 78.8 88.1 67.4 79.4
Fusing ju2022fusing 87.7 86.1 15.5 27.2 40.4 36.5
LNP bi2023detecting 63.1 67.4 56.9 72.5 58.2 68.3
UnivFD ojha2023towards 89.4 88.3 44.9 61.2 53.9 60.7
AntifakePrompt chang2023antifakeprompt 91.3 92.5 89.3 91.2 90.6 91.2
SIDA huang2024sida 92.9 93.1 90.7 91.0 91.8 92.4
@@
5.4 DeepFake detection
We evaluate FakeVLM’s performance on DeepFake detection, with results on DD-VQA presented in Table 4. FakeVLM surpasses both general-purpose multimodal models and the specialized vision-language model, Common-DF, with improvements of 5.7% in Acc, 3% in F1, and 9.5 % in ROUGE_L.
@@ 5.4 DeepFake detection
Table 4: The experimental results were evaluated on the DD-VQA datasets. Common-DF-T and Common-DF-I represent text or image contrastive losses, respectively, while Common-DF-TI denotes both text and image contrastive losses.
Method DD-VQA (ECCV 2024) Acc $\uparrow\uparrow$ F1 $\uparrow\uparrow$ ROUGE_L $\uparrow\uparrow$ CSS $\uparrow\uparrow$ InternVL2-8B chen2024internvl $56.9$ $53.1$ $14.3$ $51.1$ InternVL2-40B chen2024internvl $52.5$ $57.7$ $22.2$ $54.5$ Qwen2-VL-7B wang2024qwen2 $45.7$ $58.9$ $26.6$ $56.5$ Qwen2-VL-72B wang2024qwen2 $59.5$ $57.9$ $20.5$ $56.6$ GPT-4o gpt4o $53.2$ $31.7$ $13.0$ $42.7$ Common-DF-T fu2024commonsenset2ichallengetexttoimagegeneration $83.7$ $87.6$ $57.7$ - Common-DF-I fu2024commonsenset2ichallengetexttoimagegeneration $84.9$ $88.4$ $58.8$ - Common-DF-TI fu2024commonsenset2ichallengetexttoimagegeneration $87.5$ $90.1$ $60.9$ - FakeVLM 93.2 93.1 70.4 86.6
Table 5 shows FakeVLM’s evaluation on the FF++ DeepFake detection dataset, including sub-categories like DeepFakes, Face2Face, FaceSwap, and NeuralTextures. The experimental results demonstrate that FakeVLM continues to exhibit strong performance in these tasks, comparable to or even surpassing leading deepfake expert models. It not only leads in overall detection but also exhibits balanced performance across categories, avoiding overfitting.
Table 5: Performance evaluation of FakeVLM on the FF++ DeepFake detection dataset. The results highlight FakeVLM’s robust detection performance, on par with or outperforming top expert models.
@@
$F$-DF $F$-F2F $F$-FS $F$-NT Average
FWA li2018exposing 92.1 90.0 88.4 81.2 87.7
Face X-ray li2020facexraygeneralface 97.9 98.7 98.7 92.9 95.9
SRM lee2019srmstylebasedrecalibration 97.3 97.0 97.4 93.0 95.8
CDFA lin2024fake 99.9 86.9 93.3 80.7 90.2